Evaluate:


Lim↓ n→∞ [ n/1+n² + n/4+ n² +n/9+ n² +...... +n/2n²]

1
Expert's answer
2022-03-15T08:54:23-0400

ANSWER : "\frac{\pi}{4}"

EXPLANATION

We transform the sum "\frac{n}{1+n^2}+\frac{n}{4+n^2}+\frac{n}{9+n^2}+ \cdots+\frac{n}{2n^2}" as follows :"\frac{1 }{n}\left ( \frac{1}{\left (\frac{1}{n} \right ) ^{2} +1 }+\frac{1}{\left (\frac{2}{n} \right ) ^{2} +1 }+\cdots+\frac{1}{\left (\frac{n}{n} \right ) ^{2} +1 } \right )" = "\sum _{k=1}^{n}\frac{1}{n}\cdot \frac{1}{\left (\frac{k}{n} \right ) ^{2} +1 }." The resulting sum is the Riemann integral sum for the function "f(x)=\frac{1}{1+x^2}" on the segment [0,1] divided into "n" equal parts by points "x_{k}=\frac{k}{n}\left ( k=0,\cdots\,n \right )" . Since "x_k-x_{k-1}=\frac{1}{n}" for "k=1,\cdots,\,n" then "\sum _{k=1}^{n}\frac{1}{n}\cdot \frac{1}{\left (\frac{k}{n} \right ) ^{2} +1 } =\sum _{k=1}^{n}\ \left ( x_k-x_{k-1} \right ) \cdot \ f(x_k)" .

The function "f" is integrable, because "f" is continuous. Therefore "\lim _{n\rightarrow\infty}\sum _{k=1}^{n}\frac{1}{n}\cdot \frac{1}{\left (\frac{k}{n} \right ) ^{2} +1 } =\lim_{n\rightarrow\infty}\sum _{k=1}^{n}\ \left ( x_k-x_{k-1} \right ) \cdot \ f(x_k)=\int_{0}^{1}f(x)dx" .

"\int_ {0}^{1}\frac{1}{1+x ^{2}}dx=\left [ \arctan 1-\arctan 0 \right ] =\frac{ \pi }{4}" . Hence "\lim _{n\rightarrow\infty}\sum _{k=1}^{n}\frac{1}{n}\cdot \frac{1}{\left (\frac{k}{n} \right ) ^{2} +1 } = \frac{\pi}{4}" .